# Continual Adversarial Defense

### v4

@@ Introduction
The initial defense model $f_{0}$ will be trained in the Embedding Reservation Section. After this, we freeze the feature extractor $\varphi$ and fine-tune the expanded classifier part for the new attack $A_{i}$ using the few-shot budget $\mathcal{A}_{\mathrm{train}}^{i}$ in the Prototype Augmentation Section.
For evaluation, we are just focusing on which class the image is placed in rather than being threatened by which attack. Thus the prediction for an instance $\mathbf{x}$ is:
$y_{\mathrm{pred}}=(\arg\max f_{i}(\mathbf{x}))\%N$ (2)
Embedding Reservation for Few-Shot Adaptation
A problem brought by the few-shot budget in Principle 2 is over-fitting which is common in few-shot learning (Snell, Swersky, and Zemel 2017). To tackle this issue, we choose to reserve embedding space for future attacks by generating and assigning virtual prototypes, squeezing the embedding space of previous attacks.
First, before training the defense model $f_{0}$ at the 00-th stage, several virtual prototypes $P_{v}=[\mathbf{p}_{v}^{1},\cdots,\mathbf{p}_{v}^{V}]\in\mathbb{R}^{d\times V}$ are pre-assign in the classifier and treated as ”virtual classes” (Zhang et al. 2021), where $V=N\times T$ is the number of virtual classes, i.e. the reserved classes for future attacks. Therefore the output of current model is $f_{0}(\mathbf{x})=[W^{0},P_{v}]^{\top}\varphi(\mathbf{x})$.
Second, virtual instances are constructed by manifold mixup (Verma et al. 2019):
$\mathbf{z}=h_{2}(\lambda h_{1}(\mathbf{x}_{\mathrm{adv},r}^{0})+(1-\lambda)h_{1}(\mathbf{x}_{\mathrm{adv},s}^{0})),$ (3)
where $\mathbf{x}_{\mathrm{adv},r}$ and $\mathbf{x}_{\mathrm{adv},s}$ belong to different classes $r$ and $s$, and $\lambda\sim\mathrm{B}(\alpha,\beta)$ is a trade-off parameter the same as (Zhou et al. 2022). $h_{1}$ and $h_{2}$ are decoupled hidden layers of feature extractor i.e. $\varphi(\mathbf{x})=h_{2}\circ h_{1}(\mathbf{x})$.
@@
where $\hat{y}=\mathrm{argmax}_{j}\mathbf{p}_{v}^{j\top}\varphi(\mathbf{x}_{\mathrm{adv}}^{0})+N$ is the virtual class with maximum logit, acting as the pseudo label. $\hat{\hat{y}}=\mathrm{argmax}_{k}\mathbf{w}_{k}^{0\top}\mathbf{z}$ is the pseudo label among current known classes. $\gamma$ is a trade-off parameter, $F_{\mathrm{ce}}$ represents the standard cross-entropy loss (Zhang and Sabuncu 2018), and function $\mathsf{Mask}(\cdot)$ masks out the logit corresponding to the ground-truth:
@@
In Eq. Embedding Reservation for Few-Shot Adaptation, it pushes an instance away from the reserved virtual classes (item 1) while towards known classes (item 2), and forces a virtual instance away from known class (item 3) while towards virtual class (item 4). Trained with $\mathcal{L}_{0}$, the embedding of initial benign classes will be more compact, and the embedding spaces for virtual classes will be reserved (Zhou et al. 2022). The reserved space allows the defense model to be adapted more easily in the future and alleviates overfitting brought by few-shot budget.
Prototype Augmentation for Memory Efficiency
@@
When learning new classes, the decision boundary for previous classes can be dramatically changed, and the unified classifier is severely biased (Zhu et al. 2021). Many CL methods store a fraction of old data to jointly train the model with current data to address this issue (Rebuffi et al. 2017). However, preserving old budget may also lead to memory shortage. For memory efficiency, we adopt prototype augmentation (Zhu et al. 2022) to maintain the decision boundary of previous stages, without saving any budget. We memorize one prototype in the deep feature space for each class with each attack, and over-sample (i.e., $U_{B}$) prototypes $\mathbf{p}_{B}$ and ground-truths $y_{B}$ to the batch size, achieving the calibration of the classifier:
$\mathbf{p}_{B}=U_{B}(P_{e}),\mathcal{L}_{\mathrm{proto}}=F_{\mathrm{ce}}(\mathbf{p}_{B},y_{B})$ (6)
where $P_{e}=\{\mathbf{p}_{e}^{j}\}_{j=0}^{N\times i}$ is the set of class-wise average embedding (i.e., prototype cache) generated by feature extractor $f_{e}^{i}$ following ProtoNet (Snell, Swersky, and Zemel 2017):
$\mathbf{p}_{e}^{j}=\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}(y_{k}^{i}=j)\varphi(\mathbf{x}_{\mathrm{adv},k}^{i})$ (7)
where $\mathbb{I}(\cdot)$ is the indicator function, and $K$ is the amount of budget for each class.
@@
$\mathcal{L}_{\mathrm{ce}}=F_{\mathrm{ce}}(f_{i}(\mathbf{x}_{\mathrm{adv}}^{i}),y^{i}),\quad\mathcal{L}_{f}=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\mathrm{proto}}$ (8)
Model Ensemble for Clean Data Classification
To maintain high classification accuracy on both clean and adversarial data mentioned in Principle 4, we propose the model ensemble as the last part of CAD. Ensemble adversarial training (EAT) (Tramèr et al. 2018) which trains a robust model using adversarial samples generated by the target model is a simple yet effective way to defend against adversarial attack under the gray-box setting while maintaining the classification performance of clean data. We extend EAT to our scenario by training a small weight estimator model $f_{p}$ to fuse the logits of models.
We adopt self-perturbation (Wang et al. 2023a) to train the weight estimator at the first stage. Agnostic to any of the attacks, a self-perturbation-trained model is able to distinguish images from benign and adversarial. The weight estimator $f_{p}$ outputs a weight vector $\mathbf{w}\in\mathbb{R}^{2}$ to ensemble the defense model and the target model:
@@
Experiment
To validate the performance of CAD against various adversarial attack methods in the dynamic scenario, we conduct extensive empirical studies comparing CAD with baselines from three research streams. For evaluation and analysis, we use two metrics in this section: (1) Classification accuracy (Acc) against each attack after adaptation for defense performance evaluation. (2) Average accuracy of all attacks occurring within and before each stage, i.e. average incremental accuracy (AIAcc) (Rebuffi et al. 2017), for assessing the knowledge retention of the model.
Datasets
Experiments are conducted on CIFAR-10 and ImageNet-100. CIFAR-10 (Krizhevsky, Hinton et al. 2009) contains 5000 images for training and 1000 images for testing in each of the 10 classes. ImageNet-100 (Tian, Krishnan, and Isola 2020) is a subset of ImageNet (Deng et al. 2009) and contains 100 classes with 1000 images for training and 100 images for testing in each class. Images from CIFAR-10 and ImageNet-100 are resized to $32\times 32$ and $224\times 224$ respectively. Before training, the above datasets endured data augmentation including horizontal flip and random crop.
Attack Algorithms
We choose PGD-$l_{\infty}$ (Madry et al. 2017) as the initial attack. After that, we select 8 adversarial attack methods under $l_{\infty}$: BIM (Kurakin, Goodfellow, and Bengio 2016), RFGSM (Tramèr et al. 2018), MIM (Dong et al. 2018), DIM (Wu et al. 2021), NIM (Lin et al. 2019), SNIM (Lin et al. 2019), VNIM and VMIM (Wang and He 2021) to compose the attack pool in defense performance evaluation. The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following (Zhang et al. 2019). As for evaluating the capability of CAD in migrating catastrophic forgetting, $l_{\infty}$ attack VMIM and SNIM, $l_{2}$ attack CW (Carlini and Wagner 2017) and DeepFool (Moosavi-Dezfooli, Fawzi, and Frossard 2016), $l_{1}$ attack EAD (Chen et al. 2018) and EADEN (Chen et al. 2018), and $l_{0}$ attack OnePixel (Su, Vargas, and Sakurai 2019) and SparseFool (Modas, Moosavi-Dezfooli, and Frossard 2019) compose the attack pool. Details of attack parameters are given in the supplementary.
Stage:Attack 0:PGD 1:SNIM 2:BIM 3:RFGSM 4:MIM 5:DIM 6:NIM 7:VNIM 8:VMIM Clean
@@
TRADES (Zhang et al. 2019) 46.5 45.5 49.6 46.2 48.0 47.7 46.7 48.2 49.5 76.2
GAIRAT (Zhang et al. 2020) 59.8 60.2 58.4 59.2 58.8 58.9 58.9 59.7 59.8 74.1
RPF (Dong and Xu 2023) 56.4 49.9 51.3 56.6 54.4 52.6 53.9 53.3 50.1 69.4
DMAT (Wang et al. 2023b) 64.3 65.1 66.2 65.9 65.4 64.4 64.1 65.8 65.6 75.9
ADP (Yoon, Hwang, and Lee 2021) 77.2 76.9 71.8 71.8 75.6 75.9 76.2 76.0 72.3 74.1
DiffPure (Nie et al. 2022) 53.2 53.0 52.4 55.8 53.9 53.4 52.6 54.7 52.3 73.6
@@
Table 1: Classification accuracy(%) against attacks($l_{\infty}$) on ImageNet-100. Volumn ”Clean” represents the standard accuracy of clean images. $\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD in case no constraints on memory and defense budget, and is considered to be the empirical upper bound marked with a gray background.
Stage:Attack 0:PGD 2:SNIM 3:VNIM 4:CW 5:DeepF. 6:EAD 7:EADEN 8:OnePixel 9:SparseF.
@@
EWC (Kirkpatrick et al. 2017) 95.5 68.2 67.0 23.7 23.6 23.7 26.9 25.9 22.2
SSRE (Zhu et al. 2022) 89.1 82.5 78.1 65.8 58.4 50.6 47.5 42.4 37.8
BEEF (Wang et al. 2022) 94.9 83.1 84.7 82.3 77.9 74.2 73.8 72.6 64.5
FACT (Zhou et al. 2022) 94.5 87.2 86.4 81.5 80.9 78.6 78.4 78.2 74.6
@@
Table 2: Continual defense performance (average incremental accuracy(%)) against attacks under different norms $l_{p=0,1,2,\infty}$ on CIFAR-10. DeepF. and SparseF. denote DeepFool and SparseFool respectively.
@@
Both the target model and defense model use ResNet (He et al. 2016) as the backbone network in our method. Specifically, we use WideResNet-28-10 for CIFAR-10 following (Grathwohl et al. 2020) and ResNet-50 for ImageNet-100 following (Laidlaw, Singla, and Feizi 2020). A 4-layer ConvNet is used as the weight estimator model. The cosine classifier (Gidaris and Komodakis 2018) is adopted as the classifier of the defense model. We utilize Torchattacks (Kim 2020) to generate adversarial images. Shots of defense budget per class is set to $K=10$. The number of stages is set to $T=8$ since there are 9 attacks (PGD is the initial attack). Parameters of Beta distribution are set to $\alpha=\beta=2$ and the trade-off parameter for $\mathcal{L}_{0}$ is set to $\gamma=0.01$ following (Zhou et al. 2022). Epochs for training and fine-tuning are set to $100$ and $4$ respectively.
@@
We first compare with adversarial training methods TRADES (Zhang et al. 2019), GAIRAT (Zhang et al. 2020), RPF (Dong and Xu 2023) and DMAT (Wang et al. 2023b), and purification methods ADP (Yoon, Hwang, and Lee 2021) and DiffPure (Nie et al. 2022) in defense performance evaluation. As for evaluating the capability of CAD in migrating catastrophic forgetting, we compare to CL methods EWC (Kirkpatrick et al. 2017), SSRE (Zhu et al. 2022), BEEF (Wang et al. 2022), and FACT (Zhou et al. 2022). Besides, we propose another variant of our method as the empirical upper bound, denoted as premium-CAD. For premium-CAD, there is no constraint of memory and no concerns about the amount of defense budget. We train additional attack-specific defense models using abundant defense budget in each stage, and ensemble them with the target model by the weight estimator. For details about the implementation of defense baseline and premium-CAD, please refer to supplementary materials.
Comparisons to Baselines
To evaluate the defense performance, we first compared CAD to defense baseline methods in defending against 9 adversarial attacks. For CAD, the defense environment is dynamic, with attacks occurring at different stages. The other baseline methods operate in a static defense environment that does not differentiate the order of attacks. As depicted in Table 1, despite the small scale of the defense budget, CAD demonstrates strong adaptability to each type of attack. When compared with the top baseline DMAT and ADP on ImageNet-100, CAD outperforms them by approximately $6\%$ for each attack. This underscores its capacity for continuous adaptation in response to emerging attacks. Another notable advantage of CAD is its ability to maintain high classification performance for clean images, achieving $88.1\%$ accuracy, while the target model stands at $89.8\%$. This can be attributed to the generic representation learned by the weight estimator for adversarial examples, enabling it to effectively differentiate between clean and adversarial images.
@@
Ablation Study
As argued, embedding reservation facilitates easier adaptation of the defense model and mitigates overfitting. Prototype augmentation preserves the decision boundary of previous stages, while model ensemble ensures robust classification performance for both clean and adversarial images.
Stage:Attack 0:PGD 1:CW 2:EAD 3:OnePixel Clean

### v5

@@ 4.1 Continual Adaptation to New Attacks
The initial defense model $f_{0}$ will be trained in Section 4.2. After this, we freeze the feature extractor $\varphi$ and fine-tune the expanded classifier part for the new attack $A_{i}$ using the few-shot budget $\mathcal{A}_{\mathrm{train}}^{i}$ in Section 4.3.
Due to the internal similarity of some adversarial attacks [33], trained with one attack, the defense model can, to some extent, generalize to other attacks, resulting in some unnecessary class expansion. Old attacks occurring in new stages also lead to unnecessary adaptation. Therefore, before fine-tuning at each stage, we perform model testing to filter out the useless data from the defense budget that can be correctly classified.
For evaluation, we focus solely on the class assignment of the image, rather than the specific attack it is subjected to. Therefore, the prediction for an instance $\mathbf{x}$ is:
$y_{\mathrm{pred}}=(\arg\max f_{i}(\mathbf{x}))\bmod{N}$ (2)
4.2 Embedding Space Reservation for Few-Shot Adaptation
A problem brought by the few-shot budget in Principle 2 is over-fitting which is common in few-shot learning [27]. To tackle this issue, we choose to reserve embedding space for future attacks by generating and assigning virtual prototypes, squeezing the embedding space of previous attacks.
First, before training the defense model $f_{0}$ at stage 00, several virtual classes $W_{v}=[\mathbf{w}_{v}^{1},\cdots,\mathbf{w}_{v}^{V}]\in\mathbb{R}^{d\times V}$ are pre-assign in the classifier [39], where $V=N\times T$ is the number of virtual classes, i.e. the reserved classes for future attacks. Therefore the output of current model is $f_{0}(\mathbf{x})=[W^{0},W_{v}]^{\top}\varphi(\mathbf{x})$. After training, $W_{v}$ will be used to initiate the new parameters of the expanded classifier.
Second, virtual instances are constructed by manifold mixup [31]:
$\mathbf{z}=h_{2}(\omega h_{1}(\mathbf{x}_{\mathrm{adv},r}^{0})+(1-\omega)h_{1}(\mathbf{x}_{\mathrm{adv},s}^{0})),$ (3)
where $\mathbf{x}_{\mathrm{adv},r}$ and $\mathbf{x}_{\mathrm{adv},s}$ belong to different classes $r$ and $s$, and $\omega\sim\mathrm{B}(\alpha,\beta)$ is a trade-off parameter the same as [43]. $h_{1}$ and $h_{2}$ are decoupled hidden layers of feature extractor i.e. $\varphi(\mathbf{x})=h_{2}(h_{1}(\mathbf{x}))$.
@@ 4.2 Embedding Space Reservation for Few-Shot Adaptation
where $\hat{y}=\mathrm{argmax}_{j}\mathbf{w}_{v}^{j\top}\varphi(\mathbf{x}_{\mathrm{adv}}^{0})+N$ is the virtual class with maximum logit, acting as the pseudo label. $\hat{\hat{y}}=\mathrm{argmax}_{k}\mathbf{w}_{k}^{0\top}\mathbf{z}$ is the pseudo label among current known classes. $\gamma$ is a trade-off parameter, $F_{\mathrm{ce}}$ represents the standard cross-entropy loss [42], and function $\mathsf{Mask}(\cdot)$ masks out the logit corresponding to the ground-truth:
@@
In Eq. 4.2, it pushes instance $\mathbf{x}_{\mathrm{adv}}^{0}$ away from the reserved virtual classes (item 1) while towards known classes (item 2), and forces virtual instance $\mathbf{z}$ away from known class (item 3) while towards virtual class (item 4). Trained with $\mathcal{L}_{0}$, the embedding of initial benign classes will be more compact, and the embedding spaces for virtual classes will be reserved [43]. The reserved space allows the defense model to be adapted more easily in the future and alleviates overfitting brought by few-shot budget.
4.3 Prototype Augmentation for Memory Efficiency
@@
When learning new classes, the decision boundary for previous classes can be dramatically changed, and the unified classifier is severely biased [45]. Many CL methods store a fraction of old data to jointly train the model with current data to address this issue [25]. However, preserving old budget may also lead to memory shortage. For memory efficiency, we adopt prototype augmentation [46] to maintain the decision boundary of previous stages, without saving any budget. We memorize one prototype in the deep feature space for each class with each attack, and over-sample (i.e., $U_{B}$) prototypes $\mathbf{p}_{B}$ and ground-truths $y_{B}$ to the batch size, achieving the calibration of the classifier:
$\mathbf{p}_{B}=U_{B}(P_{e}),\quad\mathcal{L}_{\mathrm{proto}}=F_{\mathrm{ce}}(\mathbf{p}_{B},y_{B})$ (6)
where $P_{e}=\{\mathbf{p}_{e}^{j}\}_{j=0}^{N\times i}$ is the set of class-wise average embeddings [27], i.e., the prototype cache. After each adaptation, prototypes of the current attack are added into $P_{e}$:
$\mathbf{p}_{e}^{j}=\frac{1}{K}\textstyle{\sum_{k=1}^{|\mathcal{A}_{\mathrm{train}}^{i}|}}\mathbb{I}(y_{k}^{i}=j)\varphi(\mathbf{x}_{\mathrm{adv},k}^{i})$ (7)
where $\mathbb{I}(\cdot)$ is the indicator function, $K$ the amount of budget for each class, and $(\mathbf{x}_{\mathrm{adv},k}^{i},y_{k}^{i})$ the $k$-th sample in $\mathcal{A}_{\mathrm{train}}^{i}$.
@@
$\mathcal{L}_{\mathrm{ce}}=F_{\mathrm{ce}}(f_{i}(\mathbf{x}_{\mathrm{adv}}^{i}),y^{i}),\quad\mathcal{L}_{f}=\mathcal{L}_{\mathrm{ce}}+\lambda\mathcal{L}_{\mathrm{proto}}$ (8)
4.4 Model Ensemble for Clean Data Classification
To maintain high classification accuracy on both clean and adversarial data mentioned in Principle 4, we propose the model ensemble as the last part of CAD. Ensemble adversarial training (EAT) [30] which trains a robust model using adversarial samples generated by the target model is a simple yet effective way to defend against adversarial attack under the gray-box setting while maintaining the classification performance of clean data. We extend EAT to our scenario by training a small weight estimator model $f_{p}$ to fuse the logits of models.
We adopt self-perturbation [33] to train the weight estimator at the first stage. Agnostic to any of the attacks, a self-perturbation-trained model is able to distinguish images from benign and adversarial. The weight estimator $f_{p}$ outputs a weight vector $\mathbf{w}\in\mathbb{R}^{2}$ to ensemble the defense model and the target model:
@@ 4.4 Model Ensemble for Clean Data Classification
5 Experiment
To validate the performance of CAD against various adversarial attack methods in the dynamic scenario, we conduct extensive empirical studies comparing CAD with baselines from three research streams. For evaluation and analysis, we use two metrics in this section: (1) Classification accuracy (Acc) after each adaptation for defense performance evaluation. (2) Average accuracy of all attacks occurring within and before each stage, i.e. average incremental accuracy (AIAcc) [25], to assess the model’s knowledge retention across all attacks within and before each stage.
Method Classification accuracy(%) against $l_{\infty}$ attacks at each stage. Clean
0:PGD 1:SNIM 2:BIM 3:RFGSM 4:MIM 5:DIM 6:NIM 7:VNIM 8:VMIM
@@ 5 Experiment
TRADES [40] 46.5 45.5 49.6 46.2 48.0 47.7 46.7 48.2 49.5 76.2
CLC [2] 48.4 48.2 49.8 47.3 49.6 48.8 47.9 49.1 50.5 77.5
GAIRAT [41] 59.8 60.2 58.4 59.2 58.8 58.9 58.9 59.7 59.8 74.1
DMAT [35] 64.3 65.1 66.2 65.9 65.4 64.4 64.1 65.8 65.6 75.9
MeanSparse [1] 69.4 69.3 70.5 70.1 71.3 71.8 70.0 71.6 71.4 78.8
ADP [38] 77.2 76.9 71.8 71.8 75.6 75.9 76.2 76.0 72.3 74.1
DiffPure [24] 53.2 53.0 52.4 55.8 53.9 53.4 52.6 54.7 52.3 73.6
@@
Table 1: Adversarial defense performance on ImageNet-100. Volumn ”Clean” denotes standard accuracy on clean images. $\mathrm{\textbf{CAD}}^{\dagger}$ represents premium-CAD without memory or defense budget constraints, serving as the empirical upper bound (gray background).
Method Average incremental accuracy(%) against $l_{p=0,1,2,\infty}$ attacks at each stage.
0:PGD 1:SNIM 2:VNIM 3:CW 4:DeepF. 5:EAD 6:EADEN 7:OnePixel 8:SparseF.
@@
EWC [14] 95.5 68.2 67.0 23.7 23.6 23.7 26.9 25.9 22.2
SSRE [46] 89.1 82.5 78.1 65.8 58.4 50.6 47.5 42.4 37.8
BEEF [32] 94.9 83.1 84.7 82.3 77.9 74.2 73.8 72.6 64.5
FACT [43] 94.5 87.2 86.4 81.5 80.9 78.6 78.4 78.2 74.6
@@
Table 2: Continual defense performance on CIFAR-10. DeepF. and SparseF. denote DeepFool and SparseFool respectively.
Datasets
Experiments are conducted on CIFAR-10 and ImageNet-100. CIFAR-10 [15] contains 5000 images for training and 1000 images for testing in each of the 10 classes. ImageNet-100 [29] is a subset of ImageNet [4] and contains 100 classes with 1000 images for training and 100 images for testing in each class. Images from CIFAR-10 and ImageNet-100 are resized to $32\times 32$ and $224\times 224$ respectively. Before training, the above datasets endured data augmentation including horizontal flip and random crop.
Attacks
We choose PGD-$l_{\infty}$ [23] as the initial attack. After that, we select 8 adversarial attack methods under $l_{\infty}$: SNIM [20], BIM [16], RFGSM [30], MIM [6], DIM [36], NIM [20], VNIM and VMIM [34] to compose the attack pool in defense performance evaluation. The perturbation magnitude is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following [40]. To evaluate CAD’s effectiveness in mitigating catastrophic forgetting, $l_{\infty}$ attacks (VMIM and SNIM), $l_{2}$ attacks (CW and DeepFool), $l_{1}$ attacks (EAD and EADEN), and $l_{0}$ attacks (OnePixel and SparseFool) compose the attack pool. Details of attack parameters are given in the supplementary materials.
@@
Both the target and defense models use ResNet [9] as the backbone: WideResNet-28-10 for CIFAR-10 following [8] and ResNet-50 for ImageNet-100 following [17]. A 4-layer ConvNet is used as the weight estimator model. The cosine classifier [7] is adopted as the classifier of the defense model. We utilize Torchattacks [13] to generate adversarial images. The defense budget per class is set to $K=10$, with $T=8$ stages since there are 9 attacks (PGD as the initial attack). Following [43], Beta distribution parameters are $\alpha=\beta=2$, and trade-off parameters are $\gamma=0.01$. Scaling hyperparameter of prototype augmentation is set to $\lambda=0.5$. Epochs for training and fine-tuning are set to $100$ and $4$ respectively.
@@
We compare CAD with robust models TRADES [40], CLC [2], GAIRAT [41], DMAT [35], and MeanSparse [1], and purification methods ADP [38], DiffPure [24] in defense performance. Robust models are evaluated in a black-box setting, while purification methods and CAD are tested in a gray-box setting. To assess CAD’s ability to mitigate catastrophic forgetting, we compare it with CL methods EWC [14], SSRE [46], BEEF [32], and FACT [43] in a gray-box setting. Additionally, we propose a variant of CAD, premium-CAD, as the empirical upper bound, which has no memory or defense budget constraints. In premium-CAD, attack-specific defense models are trained with abundant defense budget and ensembled with the target model via a weight estimator. Details of the baseline and premium-CAD implementation can be found in the supplementary materials.
@@
5.1 Comparisons to Baselines
To evaluate the defense performance, we first compared CAD to defense baseline methods in defending against 9 adversarial attacks. For CAD, the defense environment is dynamic, with attacks occurring at different stages. The other baseline methods operate in a static defense environment that does not differentiate the order of attacks. As depicted in Table 1, despite the small scale of the defense budget, CAD demonstrates strong adaptability to each type of attack. When compared with the top baseline MeanSparse [1] and ADP [38] on ImageNet-100, CAD outperforms them by approximately $6\%$ for each attack. This underscores its capacity for continuous adaptation in response to emerging attacks. Another notable advantage of CAD is its ability to maintain high classification performance for clean images, achieving $88.1\%$ accuracy, while the target model stands at $89.8\%$. This can be attributed to the generic representation learned by the weight estimator for adversarial examples, enabling it to effectively differentiate between clean and adversarial images.
Ablation 0:PGD 1:CW 2:EAD 3:OnePixel Clean